# CUB: Benchmarking Context Utilisation Techniques for Language Models

### v1

@@ B.1 CounterFact
Samples from the CounterFact dataset can be found in Table 8. The relations covered by the dataset are capital of (80%), country of origin (9%), location of formation (9%), field of work (1%) and country of citizenship (1%).
@@
We evaluate all Regular LMs on the samples from CUB CounterFact without context. The results can be found in Table 9. We observe rates above 70% for all models. As expected, the highest memorisation rate is found for Pythia. The lowest is found for GPT-2 XL, which can be expected as the model is quite small and old.
@@
Following the same approach as previous work, no specific prompt template was used for the LMs evaluated on CounterFact. The LMs were evaluated in a simple sentence completion format as shown in Table 8.
@@
We retain all samples from the development set of NQ [^5] https://console.cloud.google.com/storage/browser/natural_questions/v1.0/dev for which a short answer of fewer than five tokens is identified in the raw HTML of the corresponding Wikipedia pages. Samples from the NQ dataset can be found in Table 10.
@@
A manual inspection of 200 samples found the method reliable for producing adequate conflicting contexts with an accuracy of 90% (11 samples corresponded to poor formatting, 4 were too similar to gold, and 4 were dropped due to data formatting issues or the LLM being unable to generate a substitute answer not already found in the context). In addition, we inspect the CUB results to ascertain the quality of the conflicting context sampling, see Appendix A.
@@ B.3 DRUID
Samples from the DRUID dataset can be found in Table 11. The evidence stance and fact-check verdict distributions per context type can be found in Tables 12 and 13.
@@
FACEBOOK POST WikiLeaks has published the 1st list of black money holders in Swiss banks. False (See attached file: List of Black Money Holders from Wiki Conflicting
Irish Congress of Trade Unions (ICTU) One in five school staff in Northern Ireland are assaulted at least once a week. False Finnegan, who died in January 2002, had also abused boys at St. Colman’s College, a prestigious Catholic boys’ secondary school in Newry, Northern Ireland. He taught there from 1967 to 1971 and again from 1973 to 1976, when he was appointed president of the school. He served in that post until 1987. […] Admitted on October 9, 2014 to sample charges of indecently assaulting four boys as young as 10 at St Mary’s CBS primary school in Mullingar between 1984 and 1987. Jailed for two years at Mullingar Circuit Court sitting in Tullamore. This concluded a ten-year investigation by detectives in Mullingar. […] When Smyth returned to Kilnacrott in 1983, he again began abusing children in Belfast, including the girl who, on February 23, 1990, would meet with a social worker at the Catholic Family Welfare Society in Belfast and start all the Smyth revelations. Irrelevant
Table 11: DRUID samples and corresponding context types.
@@ Appendix C CCU metric
$\mathrm{CCU}(t)=\begin{cases}\frac{P_{M}(t|Q,C)-P_{M}(t|Q)}{1-P_{M}(t|Q)}\\ \text{if }\scriptstyle P_{M}(t|Q,C)\geq P_{M}(t|Q),\vspace{0.3cm}\\ \frac{P_{M}(t|Q,C)-P_{M}(t|Q)}{P_{M}(t|Q)}\\ \text{otherwise.}\end{cases}$ (1)
For relevant contexts $C$ we record $\mathrm{CCU}(t_{C})$, i.e. the scores for the token promoted by the context. For irrelevant contexts we record the $\mathrm{CCU}(t_{M})$, i.e. the scores for the top token predicted by the model when prompted without context (memory). The range of $\mathrm{CCU}$ is $[-1,1]$, for which a value of $-1$ denotes that the model goes completely against the context when the context is relevant or against its memory when the context is irrelevant, and vice versa for $\mathrm{CCU}$ values of 1. We report the averaged $\mathrm{CCU}$ per context type.
@@ D.1 Prompting
The tuned prompt found for each model and dataset can be found in Table 14. Different sets of prompts were experimented with depending on dataset and model type. A set of 11 to 12 prompts were produced for each of CounterFact, NQ and DRUID for the three different model types (causal LM, instruction-tuned LMs and Command A), respectively. Prompts with the same number are similar to each other across model types (e.g. Prompt #2 for Qwen2.5 on DRUID is similar to Prompt #2 for instruction-tuned Qwen2.5 on DRUID). Prompt sets across different datasets are dissimilar as they are adapted to align the instructions and few-shot examples to the given dataset. Prompt sets across different model types for the same dataset are dissimilar as small tweaks need to be applied for the instruction-tuned models that work less well in a purely causal language modelling setting, and for Command A that is a chat-based model. All prompts will be possible to view in the code repository of the paper.
@@
Table 14: The tuned prompts for each LM. default denotes that the original prompt template (seen in Appendix B) worked best. “-I” denotes instruction-tuned model versions. The source of the prompt is indicated in parenthesis.
@@ D.2 PH3
The tuned attention head configurations for PH3 can be found in Table 15. The head configurations are grouped by the top number of identified attention heads to consider and to what extent we allow mixing between context and memory heads. E.g. #25 all denotes all top-25 context and memory heads detected, #3 memory denotes the top-3 memory heads, allowing for overlap with context heads, and #1 only memory denotes memory heads detected without overlap with context heads when considering the top-1 context and memory heads.
@@ Appendix E Implementation Details of Fine-tuning
We fine-tune the LMs with a learning rate of 5e-5, [^9] Experiments with other learning rates yielded insignificant changes in performance on the validation set. using warm-up. To avoid overfitting, we use early stopping based on the loss on the validation set. For QA datasets, we use the train split from SQuAD 2.0 Rajpurkar et al. (2018), and TriviaQA Joshi et al. (2017). For a FC dataset, we take the train split from AVeriTeC Schlichtkrull et al. (2023). For a sentence completion dataset, we take the static partition of the DYNAMICQA Marjanovic et al. (2024). We only create counterfactual training examples with DYNAMICQA dataset. The detailed statistics for mixing the selected datasets can be found in Table 17.
@@ Appendix F Additional Details of Multi-agent
7:     return $\text{LM}(a\mid q)$triangleright Answer w/o $c$
@@
13:     return $a_{c}$triangleright Answer w/ $c$
@@
18:return $\text{LM}(a\mid q,c,a_{c},f_{\text{faith}})$triangleright Self-Refined
@@
Appendix G Input Features
We detect the input features described in Section 5.2 as follows:
@@ Appendix G Input Features
Appendix H Computational Resources

### v2

@@ Appendix B Data Collection
We collect three datasets for CUB: CounterFact, NQ, and DRUID. For each dataset, we carefully construct validation and test splits to enable fair and unified hyperparameter tuning for each CMT, as well as unbiased final evaluation metrics. The following subsections describe the construction and characteristics of each dataset in detail.
@@ B.1 CounterFact
Samples from the CounterFact dataset can be found in Table˜8. The relations covered by the dataset are capital of (80%), country of origin (9%), location of formation (9%), field of work (1%) and country of citizenship (1%).
@@
We evaluate all Regular LMs on the samples from CUB CounterFact without context. The results can be found in Table˜9. We observe rates above 70% for all models. As expected, the highest memorisation rate is found for Pythia. The lowest is found for GPT-2 XL, which can be expected as the model is quite small and old.
@@
Following the same approach as previous work, no specific prompt template was used for the LMs evaluated on CounterFact. The LMs were evaluated in a simple sentence completion format as shown in Table˜8.
@@
We retain all samples from the development set of NQ [^5] https://console.cloud.google.com/storage/browser/natural_questions/v1.0/dev for which a short answer of fewer than five tokens is identified in the raw HTML of the corresponding Wikipedia pages. Samples from the NQ dataset can be found in Table˜10.
@@
We create conflicting contexts that promote a different answer simply by taking the gold context and substituting the gold answer in the context. The substitute answer is sampled to yield coherent conflicting contexts, and to have a different meaning compared to the gold answer.
@@
A manual inspection of 200 samples found the method reliable for producing adequate conflicting contexts with an accuracy of 90% (11 samples corresponded to poor formatting, 4 were too similar to gold, and 4 were dropped due to data formatting issues or the LLM being unable to generate a substitute answer not already found in the context). In addition, we inspect the CUB results to ascertain the quality of the conflicting context sampling, see Appendix˜A.
@@ B.3 DRUID
Samples from the DRUID dataset can be found in Table˜11. The evidence stance and fact-check verdict distributions per context type can be found in Tables˜12 and 13.
@@
FACEBOOK POST WikiLeaks has published the 1st list of black money holders in Swiss banks. False (See attached file: List of Black Money Holders from Wiki Conflict.
Irish Congress of Trade Unions (ICTU) One in five school staff in Northern Ireland are assaulted at least once a week. False Finnegan, who died in January 2002, had also abused boys at St. Colman’s College, a prestigious Catholic boys’ secondary school in Newry, Northern Ireland. He taught there from 1967 to 1971 and again from 1973 to 1976, when he was appointed president of the school. He served in that post until 1987. […] Admitted on October 9, 2014 to sample charges of indecently assaulting four boys as young as 10 at St Mary’s CBS primary school in Mullingar between 1984 and 1987. Jailed for two years at Mullingar Circuit Court sitting in Tullamore. This concluded a ten-year investigation by detectives in Mullingar. […] When Smyth returned to Kilnacrott in 1983, he again began abusing children in Belfast, including the girl who, on February 23, 1990, would meet with a social worker at the Catholic Family Welfare Society in Belfast and start all the Smyth revelations. Irrel.
Table 11: DRUID samples and corresponding context types. ‘Conflict.’ and ‘Irrel.’ denote conflicting and irrelevant context types, respectively.
@@
Differently from CounterFact and NQ, no context synthesis is necessary for the DRUID samples as they, by virtue of utilising naturally occurring samples from a RAG pipeline, already contain samples representative of gold, conflicting and irrelevant contexts.
@@ Appendix C CCU metric
$\mathrm{CCU}(t)=\begin{cases}\frac{P_{M}(t|Q,C)-P_{M}(t|Q)}{1-P_{M}(t|Q)}\\ \text{if }\scriptstyle P_{M}(t|Q,C)\geq P_{M}(t|Q),\vskip 8.5359pt\\ \frac{P_{M}(t|Q,C)-P_{M}(t|Q)}{P_{M}(t|Q)}\\ \text{otherwise.}\end{cases}$ (1)
For relevant contexts $C$ we record $\mathrm{CCU}(t_{C})$, i.e. the scores for the token promoted by the context. For irrelevant contexts we record the $\mathrm{CCU}(t_{M})$, i.e. the scores for the top token predicted by the model when prompted without context (memory). The range of $\mathrm{CCU}$ is [1,1][-1,1], for which a value of 1-1 denotes that the model goes completely against the context when the context is relevant or against its memory when the context is irrelevant, and vice versa for $\mathrm{CCU}$ values of 1. We report the averaged $\mathrm{CCU}$ per context type.
@@ D.1 Prompting
The tuned prompt found for each model and dataset can be found in Table˜14. Different sets of prompts were experimented with depending on dataset and model type. A set of 11 to 12 prompts were produced for each of CounterFact, NQ and DRUID for the three different model types (causal LM, instruction-tuned LMs and Command A), respectively. Prompts with the same number are similar to each other across model types (e.g. Prompt #2 for Qwen2.5 on DRUID is similar to Prompt #2 for instruction-tuned Qwen2.5 on DRUID). Prompt sets across different datasets are dissimilar as they are adapted to align the instructions and few-shot examples to the given dataset. Prompt sets across different model types for the same dataset are dissimilar as small tweaks need to be applied for the instruction-tuned models that work less well in a purely causal language modelling setting, and for Command A that is a chat-based model. All prompts will be possible to view in the code repository of the paper.
@@
Table 14: The tuned prompts for each LM. default denotes that the original prompt template (seen in Appendix˜B) worked best. “-I” denotes instruction-tuned model versions. The source of the prompt is indicated in parenthesis.
@@ D.2 PH3
The tuned attention head configurations for PH3 can be found in Table˜15. The head configurations are grouped by the top number of identified attention heads to consider and to what extent we allow mixing between context and memory heads. E.g. #25 all denotes all top-25 context and memory heads detected, #3 memory denotes the top-3 memory heads, allowing for overlap with context heads, and #1 only memory denotes memory heads detected without overlap with context heads when considering the top-1 context and memory heads.
@@ Appendix E Implementation Details of Fine-tuning
To align the domain with our evaluation data, we curate the fine-tuning data with two QA datasets Joshi et al. (2017); Rajpurkar et al. (2018), one FC dataset Schlichtkrull et al. (2023), and one sentence completion dataset Marjanovic et al. (2024). Before fine-tuning each LM, we elicit its parametric answers by querying without contexts. We then select the questions that the LM answered correctly and pair them with irrelevant and empty contexts. The fine-tuning data thus contains contexts that can be irrelevant, counterfactual, or empty. During fine-tuning, we train the LM to generate answers aligned with the provided context. When the context is irrelevant, we train the LM to be robust, i.e. ignore the context and output its parametric answer. Due to the computational costs associated with fine-tuning billion-sized LMs, we use the Low-Rank Adaptation method Hu et al. (2021).
The LMs are fine-tuned with a learning rate of 5e-5, [^9] Experiments with other learning rates yielded insignificant changes in performance on the validation set. using warm-up. To avoid overfitting, we use early stopping based on the loss on the validation set. For QA datasets, we use the train split from SQuAD 2.0 Rajpurkar et al. (2018), and TriviaQA Joshi et al. (2017). For a FC dataset, we take the train split from AVeriTeC Schlichtkrull et al. (2023). For a sentence completion dataset, we take the static partition of the DYNAMICQA Marjanovic et al. (2024). We only create counterfactual training examples with DYNAMICQA dataset. The detailed statistics for mixing the selected datasets can be found in Table˜17.
@@ Appendix F Additional Details of Multi-agent
7:  return $\text{LM}(a\mid q)$ $\triangleright$ Answer w/o $c$
@@
13:  return $a_{c}$ $\triangleright$ Answer w/ $c$
@@
18:return $\text{LM}(a\mid q,c,a_{c},f_{\text{faith}})$ $\triangleright$ Self-Refined
As illustrated in the algorithm and Figure 5, we first assess relevance using the relevance agent to determine whether the provided context should be used. Then, the faithfulness agent provides feedback on the model response that was generated with context. If the feedback indicates that the initial answer is unfaithful, the model generates a self-refined answer based on that feedback. Given that these tasks require instruction-following capabilities, we restrict our evaluation to instruction-tuned or chat LMs.
Figure 5: Overview of the multi-agent approach.
@@
Appendix G Additional Details on PH3
The PH3 method is implemented in two steps: 1) identification of attention heads responsible for context or memory reliance via path patching and 2) pruning the identified attention heads for increased memory or context usage. To identify attention heads, we use the CounterFact datasets with samples that elicit exact fact recall in each studied model (Saynova et al., 2025). For the evaluation on our studied datasets, we tune the number of heads to prune on the validation splits of each evaluation dataset, similarly to the approach by Jin et al. (2024). The attention head configuration is tuned for each mode (PH3 +memory and PH3 +context, respectively).
Appendix H Input Features
We detect the input features described in Section˜5.2 as follows:
@@ Appendix H Input Features
Appendix I Computational Resources